# LLaRA: Supercharging Robot Learning Data for Vision-Language Policy

### v1

@@ A.4 RT-2-Style baseline
For the RT-2-Style baseline, the procedure begins by normalizing each element of the action vector to a range from 0 to 1. We then quantize these values into 256 bins and map the resulting discrete values to tokens indexed from 31000 to 31255 in the tokenizer. The dataset employed is similar to inBC, but it omits the prompt for the output format, and the action is represented by the mapped special tokens. Additionally, we have prepared a version similar to $\mathcal{D}$-inBC, named $\mathcal{D}$-RT-2 Style, which takes advantage of object detection on reference images.
A.5 Real-world robot experiments
@@
In this section, we introduce the implementation details of our real-world robot experiments.
@@
The object detector for all real-world experiments is an OWLv2 [67] (ViT-base, patch size 16) running in one-shot detection mode. The object detector takes a prompt image from the same domain and detects the objects in the observation. The list of plastic toys we used as {object} (and detected) is: {duck, croissant, cupcake, blueberries, carrot, banana, grapes, donut, tomato, corn, pepper}
@@
In the GPT-4o baseline, the model is fed the same prompt as in RT-2, with an additional prompt describing the robot arm: “You are operating a robot arm at a table to complete a task specified between <task></task>. Given the current image, you need to generate a series of actions to control the robot arm to accomplish the task.”. The model’s answer is parsed to extract the pick and place points.
@@
Appendix B Expanded Experiments
In this section, we include more experiment results. All the results are collected in Tab. 12 and Tab. 13 (VIMA-0.8k), Tab. 14 and Tab. 15(VIMA-8k) and Tab. 16 and Tab. 17(VIMA-80k).
Clarification on L4 results
We adopted the testing protocol from VIMA-Bench [43], conducting evaluations across four levels of difficulty, from L1 to L4. However, during our analysis, we discovered an inconsistency in the training set: the rotation information for the robot end effector was recorded as zero for all sweeping tasks where the end effector is a spatula. Given the importance of the spatula orientation in a sweeping task and the fact that sweeping tasks constitute 25% of the evaluations at the L4 difficulty level, we concluded that our ability to accurately evaluate our method at L4 was compromised. Considering that the original VIMA model released by the authors appears to include this rotation information, we have chosen not to report the results for L4 in our study.
Ablation on action history and multi-step planning
As described in Sec. A.1.1, we enable action history and multi-step planning when generating inBC and $\mathcal{D}$-inBC. Tab. 9 shows that these designs are helpful.
Method His. Plan L1 (%) L2 (%) L3 (%)
inBC 57.3 46.2 42.9
inBC 43.5 36.9 36.7
inBC 45.8 39.6 35.8
$\mathcal{D}$-inBC + OD 63.5 51.5 50.0
$\mathcal{D}$-inBC + OD 58.5 41.5 47.1
$\mathcal{D}$-inBC + OD 53.5 39.2 37.5
Ablation on multiple image inputs
Method Mul. L1 (%) L2 (%) L3 (%)
inBC 57.3 46.2 42.9
inBC 44.6 27.7 34.2
$\mathcal{D}$-inBC + OD 63.5 51.5 50.0
$\mathcal{D}$-inBC + OD 27.7 24.6 34.2
$\mathcal{D}$-inBC + Aux (C) 64.6 58.8 49.6
$\mathcal{D}$-inBC + Aux (C) 31.2 28.5 27.1
Ablation on object detector
Figure 13: Impact of different object detectors.
We study three types of object detectors in this paper: the VLM model itself, an external object detector separately trained on the training set (the methods with a suffix OD), and an oracle from the groundtruth dataset (Oracle). Fig. 13 shows that a reliable object detector is highly beneficial, enhancing the accuracy of image-based inputs and consequently improving model performance.
Few-shot experiments using LLM / VLM
We expanded our evaluation to include the few-shot performance of existing Large Language Models (LLMs) and Vision Language Models (VLMs). Consistent with the inference settings described in Sec. A.3, we provided three additional examples from either inBC or $\mathcal{D}$-inBC for each test scenario. The results of these evaluations are detailed in Tab. 11, where we examine three variables:
Table 11: Three-shot performance of LLMs / VLMs
Output Vision Oracle Same-task L1 (%) L2 (%) L3 (%) L4 (%)
GPT-4o 1.4 1.4 0.0 N/A
GPT-4o 0.4 0.4 0.4 0.0
GPT-4o 1.4 1.8 0.5 N/A
GPT-4o 1.2 0.4 0.4 1.2
GPT-4o 5.9 4.5 4.0 N/A
GPT-4o 6.9 5.8 4.2 1.2
Llama 3 70B (4-bit) 0.0 0.0 0.0 N/A
Llama 3 70B (4-bit) 0.8 1.2 0.4 1.2
Llama 3 70B (4-bit) 0.0 1.4 0.5 N/A
Llama 3 70B (4-bit) 0.0 0.4 0.4 0.0
LlaVA-1.5-7B 0.9 1.4 0.5 N/A
LlaVA-1.5-7B 0.0 0.0 1.0 1.2
LlaVA-1.6-34B 1.4 1.4 0.5 N/A
LlaVA-1.6-34B 1.2 0.8 2.1 0.0
Appendix C Limitations
@@
First, some concepts in images still can not be easily and precisely described by language, while in many cases these features are critical for robot control.
Second, we still rely on the object detector trained on a limited number of classes which limit the generalization ability of this method.
Third, the information extracted from the dataset can be noisy, which may lead to model confusion. In Tab. 1, there are discrepancies such as the location of an object in a reference image differing from its actual location in the current image observation. At the same time, in VIMA dataset, when referring to the shape of an object without a color, the object in the reference image will appear in gray, which is a totally different color than the object with the same shape in the current observations.
@@
Table 17: Results on VIMA-80k dataset with object detector or Oracle. Ep: number of epoch; U: normalized number of model iterations. The following sizes of datasets are relative to the size of inBC. Loc.: relative size of the localization dataset; Det.: relative size of the detection dataset; Act.: relative size of the action prediction dataset; Fut.: relative size of the future prediction dataset; Spa.: relative size of the spatial relationship dataset; Temp.: relative size of the temporal relationship dataset.
@@
‘Could you detail the steps needed to transform the scene shown in the image into the second scene? The second scene is provided as a collection of object bounding boxes {scene}. The format for these bounding boxes is <b>(x, y), {w, h}</b>, where x and y represent the center coordinates, and w and h are the width and height. The coordinates should be normalized to a scale of 0 to 1, starting from the top left corner.’
‘Can you describe what actions are required to rearrange the scene shown in the image to match the second scene? The second scene is given as a set of object bounding boxes {scene}. These bounding boxes follow the format <b>(x, y), {w, h}</b>, where x and y indicate the center coordinates, and w and h represent the width and height. The coordinates should start from the top left corner and be normalized to a scale of 0 to 1.’
‘Could you list the steps necessary to modify the scene shown in the image to the second scene? The second scene is described as a collection of object bounding boxes {scene}. The bounding box format is <b>(x, y), {w, h}</b>, with x and y denoting the center coordinates, and w and h representing the width and height. The coordinates are normalized to a scale of 0 to 1, starting from the top left corner.’
‘Can you explain what needs to be done to adjust the scene shown in the image to resemble the second scene? The second scene {scene} consists of object bounding boxes provided in the format <b>(x, y), {w, h}</b>. Here, x and y represent the center coordinates, and w and h are the width and height. The coordinates should start from the top left corner and be normalized to a scale of 0 to 1.’
‘Could you outline the necessary actions to arrange the scene shown in the image into the second scene? The second scene is defined by a collection of object bounding boxes {scene}. These bounding boxes follow the format <b>(x, y), {w, h}</b>, where x and y denote the center coordinates, and w and h are the width and height. The coordinates start from the top left corner and should be normalized to a scale of 0 to 1.’
‘Can you specify what needs to be done to convert the scene shown in the image into the second scene? The second scene is provided as a series of object bounding boxes {scene}. The format for these bounding boxes is <b>(x, y), {w, h}</b>, with x and y representing the center coordinates, and w and h indicating the width and height. Coordinates should be normalized from the top left corner to a scale of 0 to 1.’
‘Could you describe the steps required to change the scene shown in the image to the second scene? The second scene is depicted as a collection of object bounding boxes {scene}. The bounding box format is <b>(x, y), {w, h}</b>, where x and y denote the center coordinates, and w and h represent the width and height. The coordinates are normalized to a scale of 0 to 1 starting from the top left corner.’
@@
‘The image shows a scene with various objects. Now you {pick and place}, what will the scene look like? List the object bounding boxes. The format for these bounding boxes is <b>(x, y), {w, h}</b>, with x and y representing the center coordinates, and w and h as the width and height. Coordinates should start from the top left corner and be normalized to a scale of 0 to 1.’
‘Displayed in the image is a scene containing multiple objects. Now you {pick and place}, what will the scene look like? Write the list of object bounding boxes. The bounding box format is <b>(x, y), {w, h}</b>, where x and y denote the center coordinates, and w and h are the width and height. The coordinates start from the top left corner and are normalized to a scale of 0 to 1.’
‘The image illustrates a scene with various objects. Now you {pick and place}, what will the scene look like? List the object bounding boxes. The bounding boxes are formatted as <b>(x, y), {w, h}</b>, where x and y indicate the center coordinates, and w and h represent the width and height. Coordinates should be normalized from the top left corner to a scale of 0 to 1.’
‘An image depicts a scene with multiple objects. Now you {pick and place}, what will the scene look like? Write the list of object bounding boxes. The bounding box format is <b>(x, y), {w, h}</b>, where x and y represent the center coordinates, and w and h denote the width and height. The coordinates should start from the top left corner and be normalized to a scale of 0 to 1.’
@@
‘Can you describe how {ego_obj} is positioned relative to {ref_obj} in this image using words such as left, right, above, below, etc.? Also, find the 2D center distance and the Euclidean center distance between them. Your output should be in this format: {example}. The coordinates are image coordinates normalized to a scale of 0 to 1 starting from the top left corner.’ ‘Could you detail the relative positions of {ego_obj} compared to {ref_obj} in this image using terms like left, right, above, below, etc.? Also, calculate the 2D center distance and the Euclidean center distance between them. Your output must be formatted as follows: {example}. The coordinates are image coordinates normalized to a scale of 0 to 1 starting from the top left corner.’ ‘Please describe the spatial relationship of {ego_obj} in comparison to {ref_obj} in this image using relative location terms such as left, right, above, below, etc. Additionally, find the 2D center distance and the Euclidean center distance between them. Your output should match this format: {example}. The coordinates are image coordinates normalized to a scale of 0 to 1 starting from the top left corner.’ ‘Can you describe the relative spatial locations of {ego_obj} compared to {ref_obj} in this image? Use relative location words like left, right, above, below, etc. Also, calculate the 2D center distance and the Euclidean center distance between them. Your output should follow this format: {example}. The coordinates are image coordinates normalized to a scale of 0 to 1 starting from the top left corner.’ ‘Could you describe the spatial locations of {ego_obj} relative to {ref_obj} in this image using words such as left, right, above, below, etc.? Additionally, find the 2D center distance and the Euclidean center distance between them. Your output must be in this format: {example}. The coordinates are image coordinates normalized to a scale of 0 to 1 starting from the top left corner.’

### v2

@@ A.4 RT-2-Style baselines
For the RT-2-Style baseline, we follow the original RT-2’s design (Brohan et al., 2023a) and adapt special tokens to present the numerical action in VIMA action space. The original action space of VIMA contains two points for pick and place. Each point contains a 2D coordinate in robot coordinates for position and a 3D orientation of the gripper. Since we only have yaw rotation in the dataset, we simplify two 3D orientations into one yaw angle. So the action space is now 5 Degrees of Freedom (DoF): 4 numerical values present the 2D pick and place locations in robot coordinates, and 1 value presents yaw rotation. In our implementation, the procedure begins by normalizing each element of the action to a range from 0 to 1. We then quantize these values into 256 bins and map the resulting discrete values to tokens indexed from 31000 to 31255 in the tokenizer. The dataset employed is similar to inBC, but it omits the prompt for the output format, and the action is represented by the mapped special tokens. Additionally, we have prepared a version similar to $\mathcal{D}$-inBC, named $\mathcal{D}$-RT-2 Style, which takes advantage of object detection on reference images. $\mathcal{D}$-RT-2 Style (I) is similar to $\mathcal{D}$-RT-2 Style. However, four special tokens present pick and place positions in two normalized 2D image coordinates for each action, and one token presents the rotation, instead of an action in VIMABench action space. It is the same action presentation generated by our LLaRA but our method responds in natural language while $\mathcal{D}$-RT-2 Style (I) responds in special tokens.
Appendix B Expanded simulation experiments
In this section, we include more experiment results. All the numerical results reported in this paper, especially in the figures, are collected in Tab. 14 and Tab. 15 (VIMA-0.8k), Tab. 16 and Tab. 17 (VIMA-8k), Tab. 18 and Tab. 19 (VIMA-80k), and Tab. 20 (full VIMA dataset).
B.1 Clarification on L4 results
We adopted the testing protocol from VIMA-Bench (Jiang et al., 2023), conducting evaluations across four levels of difficulty, from L1 to L4. However, during our analysis, we discovered an inconsistency in the training set: the rotation information for the robot end effector was recorded as zero for all sweeping tasks where the end effector is a spatula. Given the importance of the spatula orientation in a sweeping task and the fact that sweeping tasks constitute 25% of the evaluations at the L4 difficulty level, we concluded that our ability to accurately evaluate our method at L4 was compromised. Considering that the original VIMA model released by the authors appears to include this rotation information, we have chosen not to report the results for L4 in our study.
B.2 Extended comparison to VIMA and other baselines
We would first clarify the difference between our method and VIMA (Jiang et al., 2023) in terms of the inputs. VIMA takes both front and top view images from the environment while ours only takes the front view. We test the most capable model released by VIMA, which is trained on 660k expert trajectories, as well as other RT-2-Style baselines trained on the same whole 660k dataset. The results are listed in Tab. 9. Compared to VIMA (Jiang et al., 2023) and other baselines, our best model not only achieves better performance but also requires less input and is trained on only 12% of the data used in VIMA.
Table 9: Comparison to VIMA (Jiang et al., 2023) and other baselines. Our best model not only achieves better performance but also requires less input and is trained on only 12% of the data used in VIMA.
Method Config Data L1 (%) L2 (%) L3 (%)
VIMA (Jiang et al., 2023) VIMA-200M + Oracle 100% 80.7 81.9 77.9
RT-2 Style 12% 56.9 53.1 46.2
100% 65.0 59.6 42.5
$\mathcal{D}$-RT-2 Style $\mathcal{D}$-RT-2 Style + Oracle 12% 74.2 70.4 69.2
$\mathcal{D}$-RT-2 Style + Oracle 100% 86.2 88.1 75.8
$\mathcal{D}$-RT-2 Style (I) $\mathcal{D}$-RT-2 Style (I) + Oracle 12% 58.1 54.2 55.4
$\mathcal{D}$-RT-2 Style (I) + Oracle 100% 70.0 70.0 59.2
$\mathcal{D}$-inBC + Oracle 12% 87.3 85.4 82.1
LLaRA (Ours) $\mathcal{D}$-inBC + Aux (B) + Oracle 12% 90.0 88.1 79.2
$\mathcal{D}$-inBC + Aux (D) + Oracle 12% 83.8 88.1 78.8
B.3 Will baselines benefit from longer training?
The introduction of auxiliary datasets in the training set naturally increases the number of optimization steps per training epoch. In this subsection, we confirm that the performance improvement of LLaRA with auxiliary datasets is due to the data itself, rather than simply more optimization steps. To validate this, we trained several baselines without auxiliary data for longer epochs on three VIMA subsets. As shown in Fig. 11, Fig. 11, and Fig. 11, extending the training epochs does not enhance performance, indicating that the auxiliary data provides additional information that boosts model performance.
Figure 9: Model performances on VIMA-0.8k for longer epochs.
Figure 10: Model performances on VIMA-8k for longer epochs.
Figure 11: Model performances on VIMA-80k for longer epochs.
B.4 Ablation on action history and multi-step planning
As described in Appendix A.1.1, we enable action history and multi-step planning when generating inBC and $\mathcal{D}$-inBC. Tab. 10 shows that these designs are helpful.
Method His. Plan L1 (%) L2 (%) L3 (%)
inBC 57.3 46.2 42.9
inBC 43.5 36.9 36.7
inBC 45.8 39.6 35.8
$\mathcal{D}$-inBC + OD 63.5 51.5 50.0
$\mathcal{D}$-inBC + OD 58.5 41.5 47.1
$\mathcal{D}$-inBC + OD 53.5 39.2 37.5
B.5 Ablation on multiple image inputs
Method Mul. L1 (%) L2 (%) L3 (%)
inBC 57.3 46.2 42.9
inBC 44.6 27.7 34.2
$\mathcal{D}$-inBC + OD 63.5 51.5 50.0
$\mathcal{D}$-inBC + OD 27.7 24.6 34.2
$\mathcal{D}$-inBC + Aux (C) 64.6 58.8 49.6
$\mathcal{D}$-inBC + Aux (C) 31.2 28.5 27.1
B.6 Ablation on object detector
This ablation studies three types of object detectors in this paper: the VLM model itself, an external object detector separately trained on the training set (the methods with a suffix OD), and an oracle from the groundtruth dataset (Oracle). Fig. 12 shows that a reliable object detector is highly beneficial, enhancing the accuracy of image-based inputs and consequently improving model performance.
Figure 12: Impact of different object detectors.
B.7 Ablation on action prompt during inference
This ablation studies the effect of the prompt sentence used in action generation. By default, during test time we formulate an instruction similar to Tab. 5 to query the VLM. In addition, we randomly sample one sentence from Tab. 21, named action prompt, and add it to the instruction. We study the following scenarios to understand the effect of this action prompt selection:
  • random: this is the default behavior, a random action prompt will be added to each instruction;
  • no_prompt: none instruction has the action prompt;
  • prompt $p_{a}$, where $p_{a}\in[0,14]$: all the instructions in this setting use the $p_{a}^{th}$ prompt (0-index) in Tab. 21.
The results are presented in Fig. 13, where two variants of $\mathcal{D}$-inBC are trained on VIMA-80k. In general, the selection of the action prompt does not result in a significant difference in performance except that $\mathcal{D}$-inBC + Aux (D) seems to perform better on L1 and L3 without the action prompt.
Figure 13: Effect of prompt sentence for action generation.
B.8 Few-shot experiments using LLM / VLM
We expanded our evaluation to include the few-shot performance of existing Large Language Models (LLMs) and Vision Language Models (VLMs). Consistent with the inference settings described in Appendix A.3, we provided three additional examples from either inBC or $\mathcal{D}$-inBC for each test scenario. The results of these evaluations are detailed in Tab. 12, where we examine three variables:
Table 12: Three-shot performance of LLMs / VLMs
Output Vision Oracle Same-task L1 (%) L2 (%) L3 (%) L4 (%)
GPT-4o 1.4 1.4 0.0 N/A
GPT-4o 0.4 0.4 0.4 0.0
GPT-4o 1.4 1.8 0.5 N/A
GPT-4o 1.2 0.4 0.4 1.2
GPT-4o 5.9 4.5 4.0 N/A
GPT-4o 6.9 5.8 4.2 1.2
Llama 3 70B (4-bit) 0.0 0.0 0.0 N/A
Llama 3 70B (4-bit) 0.8 1.2 0.4 1.2
Llama 3 70B (4-bit) 0.0 1.4 0.5 N/A
Llama 3 70B (4-bit) 0.0 0.4 0.4 0.0
LlaVA-1.5-7B 0.9 1.4 0.5 N/A
LlaVA-1.5-7B 0.0 0.0 1.0 1.2
LlaVA-1.6-34B 1.4 1.4 0.5 N/A
LlaVA-1.6-34B 1.2 0.8 2.1 0.0
Appendix C Details on real-world robot experiments
@@
In this section, we provide more details on real-world robot experiments first introduced in Sec. 6.2.
C.1 Environment setting
We utilize an xArm7 robot arm equipped with a gripper and a Logitech C140 RGB webcam positioned above the arm to gather observations. The camera is mounted in a fixed position above and slightly in front of the arm, providing a third-person view of the scene. The captured images have a raw resolution of $640\times 480$, which are then center-cropped to $640\times 320$ to match the aspect ratio of VIMA images. The robot’s action space includes 4 DoF for pick-and-place positioning and 1 DoF for rotation, consistent with the simplified VIMA setup.
@@ C.1 Environment setting
The object detector for all real-world experiments is an OWLv2 (Minderer et al., 2024) (ViT-base, patch size 16) running in one-shot detection mode. The object detector takes a prompt image from the same domain and detects the objects in the observation. The list of plastic toys we used as {object} (and detected) is: {duck, croissant, cupcake, blueberries, carrot, banana, grapes, donut, tomato, corn, pepper}
C.2 GPT-4o baseline
@@
In the GPT-4o baseline, the model is fed the same prompt as in RT-2, with an additional prompt describing the robot arm: ‘‘You are operating a robot arm at a table to complete a task specified between <task></task>. Given the current image, you need to generate a series of actions to control the robot arm to accomplish the task.’’. The model’s answer is parsed to extract the pick and place points.
C.3 Real-world robot dataset
We further collect 1,256 real-world images from the same real-world robot setting as the in-domain data. Out of these images, 721 feature a single object placed randomly and uniformly on the table by a robotic arm, while the remaining images display multiple objects scattered across the table. We employed one-shot detection using OWLv2 (Minderer et al., 2024) to identify and generate bounding boxes for each object in the images.
@@ C.3 Real-world robot dataset
C.4 Qualitative examples on unseen real-world tasks
In Fig. 14 we qualitatively show that LLaRA can complete some unseen tasks that are not in the training dataset. In the first example, we ask LLaRA to ‘place all other objects into the large bowl’. LLaRA is able to recognize the objects to move and not to move and outputs the correct actions. In the second example, we let LLaRA ‘get the weight of the tomato and put it back’. LLaRA can put the tomato on the scale and move the tomato back to the similar place as the initial position, showing the understanding of ‘putting it back’. These examples show the potential generalization ability of LLaRA. Video demonstrations are also included in the supplementary material.
Appendix D Limitations
@@
First, some concepts in the reference images still can not be easily and precisely described by language, while in many cases these features are critical for robot control. Due to the single image limit of the current VLM, we still rely on the object detector trained on a limited number of classes to describe the additional images in the task description, which limits the generalization ability of this method. In principle, in the future, this can be addressed by a VLM trained on interleave image text data like (Bai et al., 2023; Li et al., 2024a).
Second, the information extracted from the dataset can be noisy, which may lead to model confusion. In Tab. 5, there are discrepancies such as the location of an object in a reference image differing from its actual location in the current image observation. At the same time, in VIMA dataset, when referring to the shape of an object without a color, the object in the reference image will appear in gray, which is a totally different color than the object with the same shape in the current observations.
@@
Appendix E Raw results and prompt templates
This section contains all the numerical results of VIMA experiments and all prompt templates we used in the paper.
@@ Appendix E Raw results and prompt templates
$\mathcal{D}$-RT-2 Style (I) + Oracle 1 1 1.2 0.0 0.8
$\mathcal{D}$-RT-2 Style (I) + Oracle 2 2 2.3 1.9 1.7
$\mathcal{D}$-RT-2 Style (I) + Oracle 4 4 0.8 0.4 0.8
$\mathcal{D}$-RT-2 Style (I) + Oracle 6 6 1.5 2.7 0.4
$\mathcal{D}$-RT-2 Style (I) + Oracle 8 8 0.8 1.9 0.4
@@
$\mathcal{D}$-RT-2 Style (I) + Oracle 1 1 1.5 1.2 1.2
$\mathcal{D}$-RT-2 Style (I) + Oracle 2 2 21.9 21.5 20.8
$\mathcal{D}$-RT-2 Style (I) + Oracle 4 4 25.8 23.1 19.6
$\mathcal{D}$-RT-2 Style (I) + Oracle 6 6 17.3 18.8 15.8
$\mathcal{D}$-RT-2 Style (I) + Oracle 8 8 21.9 20.0 20.4
@@
Table 19: Results on VIMA-80k dataset with Oracle. Ep: number of epoch; U: normalized number of model iterations. The following sizes of datasets are relative to the size of inBC. Loc.: relative size of the localization dataset; Det.: relative size of the detection dataset; Act.: relative size of the action prediction dataset; Fut.: relative size of the future prediction dataset; Spa.: relative size of the spatial relationship dataset; Temp.: relative size of the temporal relationship dataset.
@@
$\mathcal{D}$-RT-2 Style (I) + Oracle 1 1 57.7 54.6 57.1
$\mathcal{D}$-RT-2 Style (I) + Oracle 2 2 55.8 54.6 55.4
$\mathcal{D}$-RT-2 Style (I) + Oracle 4 4 58.1 54.2 55.4
$\mathcal{D}$-RT-2 Style (I) + Oracle 6 6 53.5 49.6 47.9
$\mathcal{D}$-RT-2 Style (I) + Oracle 8 8 53.8 50.8 48.8
@@
Table 20: Results on full VIMA dataset with Oracle. Ep: number of epoch; U: normalized number of model iterations. The following sizes of datasets are relative to the size of inBC. Loc.: relative size of the localization dataset; Det.: relative size of the detection dataset; Act.: relative size of the action prediction dataset; Fut.: relative size of the future prediction dataset; Spa.: relative size of the spatial relationship dataset; Temp.: relative size of the temporal relationship dataset.
Method Ep U Loc. Det. Act. Fut. Spa. Temp. L1 (%) L2 (%) L3 (%)
RT-2 Style 1 1 61.9 55.8 47.1
RT-2 Style 2 2 65.0 59.6 42.5
RT-2 Style 3 3 63.1 57.7 44.6
RT-2 Style 4 4 64.2 60.0 44.2
$\mathcal{D}$-RT-2 Style + Oracle 1 1 82.3 80.8 74.6
$\mathcal{D}$-RT-2 Style + Oracle 2 2 86.2 88.1 75.8
$\mathcal{D}$-RT-2 Style + Oracle 3 3 87.7 83.5 76.2
$\mathcal{D}$-RT-2 Style + Oracle 4 4 87.3 85.4 74.6
$\mathcal{D}$-RT-2 Style (I) + Oracle 1 1 68.8 68.1 59.6
$\mathcal{D}$-RT-2 Style (I) + Oracle 2 2 68.5 66.2 58.3
$\mathcal{D}$-RT-2 Style (I) + Oracle 3 3 70.0 70.0 59.2
$\mathcal{D}$-RT-2 Style (I) + Oracle 4 4 69.6 69.2 58.3
@@
Table 24: (Continued) Prompt candidates for object detection.
@@
‘Could you detail the steps needed to transform the scene shown in the image into the second scene? The second scene is provided as a collection of object bounding boxes {scene}. The format for these bounding boxes is <b>(x, y), {w, h}</b>, where x and y represent the center coordinates, and w and h are the width and height. The coordinates should be normalized to a scale of 0 to 1, starting from the top left corner.’ ‘Can you describe what actions are required to rearrange the scene shown in the image to match the second scene? The second scene is given as a set of object bounding boxes {scene}. These bounding boxes follow the format <b>(x, y), {w, h}</b>, where x and y indicate the center coordinates, and w and h represent the width and height. The coordinates should start from the top left corner and be normalized to a scale of 0 to 1.’ ‘Could you list the steps necessary to modify the scene shown in the image to the second scene? The second scene is described as a collection of object bounding boxes {scene}. The bounding box format is <b>(x, y), {w, h}</b>, with x and y denoting the center coordinates, and w and h representing the width and height. The coordinates are normalized to a scale of 0 to 1, starting from the top left corner.’ ‘Can you explain what needs to be done to adjust the scene shown in the image to resemble the second scene? The second scene {scene} consists of object bounding boxes provided in the format <b>(x, y), {w, h}</b>. Here, x and y represent the center coordinates, and w and h are the width and height. The coordinates should start from the top left corner and be normalized to a scale of 0 to 1.’ ‘Could you outline the necessary actions to arrange the scene shown in the image into the second scene? The second scene is defined by a collection of object bounding boxes {scene}. These bounding boxes follow the format <b>(x, y), {w, h}</b>, where x and y denote the center coordinates, and w and h are the width and height. The coordinates start from the top left corner and should be normalized to a scale of 0 to 1.’ ‘Can you specify what needs to be done to convert the scene shown in the image into the second scene? The second scene is provided as a series of object bounding boxes {scene}. The format for these bounding boxes is <b>(x, y), {w, h}</b>, with x and y representing the center coordinates, and w and h indicating the width and height. Coordinates should be normalized from the top left corner to a scale of 0 to 1.’ ‘Could you describe the steps required to change the scene shown in the image to the second scene? The second scene is depicted as a collection of object bounding boxes {scene}. The bounding box format is <b>(x, y), {w, h}</b>, where x and y denote the center coordinates, and w and h represent the width and height. The coordinates are normalized to a scale of 0 to 1 starting from the top left corner.’
@@
‘The image shows a scene with various objects. Now you {pick and place}, what will the scene look like? List the object bounding boxes. The format for these bounding boxes is <b>(x, y), {w, h}</b>, with x and y representing the center coordinates, and w and h as the width and height. Coordinates should start from the top left corner and be normalized to a scale of 0 to 1.’ ‘Displayed in the image is a scene containing multiple objects. Now you {pick and place}, what will the scene look like? Write the list of object bounding boxes. The bounding box format is <b>(x, y), {w, h}</b>, where x and y denote the center coordinates, and w and h are the width and height. The coordinates start from the top left corner and are normalized to a scale of 0 to 1.’ ‘The image illustrates a scene with various objects. Now you {pick and place}, what will the scene look like? List the object bounding boxes. The bounding boxes are formatted as <b>(x, y), {w, h}</b>, where x and y indicate the center coordinates, and w and h represent the width and height. Coordinates should be normalized from the top left corner to a scale of 0 to 1.’ ‘An image depicts a scene with multiple objects. Now you {pick and place}, what will the scene look like? Write the list of object bounding boxes. The bounding box format is <b>(x, y), {w, h}</b>, where x and y represent the center coordinates, and w and h denote the width and height. The coordinates should start from the top left corner and be normalized to a scale of 0 to 1.’
@@
‘Can you describe how {ego_obj} is positioned relative to {ref_obj} in this image using words such as left, right, above, below, etc.? Also, find the 2D center distance and the Euclidean center distance between them. Your output should be in this format: {example}. The coordinates are image coordinates normalized to a scale of 0 to 1 starting from the top left corner.’
‘Could you detail the relative positions of {ego_obj} compared to {ref_obj} in this image using terms like left, right, above, below, etc.? Also, calculate the 2D center distance and the Euclidean center distance between them. Your output must be formatted as follows: {example}. The coordinates are image coordinates normalized to a scale of 0 to 1 starting from the top left corner.’
‘Please describe the spatial relationship of {ego_obj} in comparison to {ref_obj} in this image using relative location terms such as left, right, above, below, etc. Additionally, find the 2D center distance and the Euclidean center distance between them. Your output should match this format: {example}. The coordinates are image coordinates normalized to a scale of 0 to 1 starting from the top left corner.’
‘Can you describe the relative spatial locations of {ego_obj} compared to {ref_obj} in this image? Use relative location words like left, right, above, below, etc. Also, calculate the 2D center distance and the Euclidean center distance between them. Your output should follow this format: {example}. The coordinates are image coordinates normalized to a scale of 0 to 1 starting from the top left corner.’
‘Could you describe the spatial locations of {ego_obj} relative to {ref_obj} in this image using words such as left, right, above, below, etc.? Additionally, find the 2D center distance and the Euclidean center distance between them. Your output must be in this format: {example}. The coordinates are image coordinates normalized to a scale of 0 to 1 starting from the top left corner.’